# Selective Prompt Anchoring for Code Generation

### v4

@@ 1 Introduction
Large language models (LLMs) have emerged as powerful programming assistants. They have demonstrated unprecedented capabilities in interpreting natural language descriptions of programming tasks and generating source code [17, 41]. Despite this great progress, LLMs still produce incorrect solutions to challenging tasks or generate code that does not fully meet user expectations [54, 37].
To improve the performance of LLMs on coding tasks, many efforts have been made to develop high-quality training data [26, 17, 56] and design new domain-specific training objectives [38, 2]. However, these approaches require tremendous computational resources. To address this challenge, various prompting methods have been explored to enhance the inference process of code LLMs, e.g., retrieval-augmented generation [12], chain-of-thoughts [24, 47], self-planning and debugging [21, 4], etc.
@@ 2 Attention Analysis of Code LLMs
We conduct an empirical study to investigate the attention dilution phenomenon in code LLMs. Following prior studies [60, 15], we obtain self-attention scores from the last layer in LLMs, which has been shown to represent the most accurate attention distribution [23, 51]. We calculate the percentage of attention on the user prompt. Calculation details are provided in Appendix B. We also experimented with an alternative gradient-based attention calculation method [42] and obtained similar results, as detailed in Appendix A.
We analyzed five code LLMs on HumanEval [3], a widely-used code generation benchmark. Figure 3 shows the shift of LLMs’ attention on the user prompt during code generation. The result shows that as more tokens are generated, models’ attention on the user prompt gradually becomes less. Consequently, the code generation process becomes increasingly influenced by tokens generated in recent time steps, rather than the user prompt. This can be problematic in two ways. First, any generation inaccuracy, such as creating an unhelpful variable in earlier steps, is likely to propagate and influence subsequent steps, as the model tends to focus more on recent code tokens. Second, when generating long and complex code sequences for tasks with many requirements in detail, the model may overlook some specifications in the later generation stages due to insufficient attention to the user prompt. Figure 3 shows that the length of incorrectly generated code is significantly longer than that of correctly generated code. We include more details in Appendix M.
@@ 3.1 Overview
We propose Selective Prompt Anchoring (Spa) to steer model attention by amplifying the contextual impact of the important tokens in the user prompt. In this work, we assume these important tokens are selected by users. Inspired by the anchoring effect [14] in psychology, we call these important tokens “anchored text”. For example, in the prompt, “Given a string, count the number of uppercase vowels”, the user may want to emphasize “uppercase vowels" to ensure that the model does not forget to define uppercase vowels (i.e., “A, E, I, O, U”) in the generated code.
@@ 3.2 Attention Steering and Prompt Anchoring
Spa performs attention steering by scaling the impact of selected tokens to the output logits . For simplicity, we demonstrate by making the entire user prompt as the anchored text. is an input embedding matrix at step , and represents a submatrix within covering the first columns (corresponding to the prompt ). They are visualized below:
@@
To construct , we retain only the first columns of that correspond to , while setting all other columns to zero. This matrix remains constant during token generation.
Conversely, is formed by zeroing out these same first columns of that correspond to , while preserving all elements in the remaining columns.
@@
encapsulates the semantics for user prompt, while encapsulates the semantics for generated code. The sum of and reconstructs the original matrix :
Suppose we want to scale the semantic impact of the matrix by a value . We refer to as anchoring strength. We use to represent the augmented embedding matrix, where the semantics of in the original embedding is scaled by times.111Scaling the semantics of by times is not equivalent to multiplying by .
@@
In this work, we focus on scaling up the impact of the user prompt to mitigate the attention dilution issue. Let represent the augmented logits calculated by model at step , where the impact of anchored text is scaled by . We can calculate the integral of the partial derivative of with respect to from 0 to .222 is differentiable for backpropagation Formally,
@@ 3.3 Augmented Logits by Approximation
where the the integral part in Equation 7 is approximated by . To calculate , 333 does not mean setting the embedding vector to zeros. Instead, it means setting to zero, which replaces the original embedding for anchored text with the mask embedding that contains no semantic information. This mask embedding vector is non-zero. Conversely, a zero embedding does not necessarily indicate the absence of semantic information. we mask the anchored text using mask embeddings. Each LLM provides special tokens reserved for text masking, which almost has no semantic impact, e.g., <unk> for Code Llama [41] and <pad> for DeepSeek-Coder [17]. Each special token corresponds to a mask embedding. By replacing embeddings of with mask embeddings, we get a masked input matrix . It ablates the semantic impact of the anchored text while ensuring that the positional encoding remains unaffected. Thus, we can get
@@
4 Experiment Setucp
@@ 4.1 Benchmarks
HumanEval [3] includes 164 Python tasks designed by OpenAI and now has become a widely-used benchmark for code generation.
MBPP [1] is another popular benchmark that includes 974 crowd-sourced Python tasks. Due to ambiguous task descriptions, the authors of MBPP created a sanitized version that included 427 tasks with clearer descriptions. We evaluate Spa on the sanitized version.
HumanEval+ and MBPP+ [31] improves the original HumanEval and MBPP benchmarks with additional test cases to cover corner cases [32].
HumanEval-X [18] extends the HumanEval benchmark to support more programming languages such as Python, Java, JavaScript, C++, and Go. It aims to evaluate the multilingual code generation abilities.
BigCodeBench [64] is a more challenging benchmark for code generation that evaluates models’ abilities to follow complex instructions and use tools, including 1,140 real-world Python tasks across 139 libraries.
@@ 4.2 Baselines
Base Models. We select six representative open-source code LLMs: CodeGen-Mono-350M [36], CodeLlama-7B [41], StarCoder2-15B [34], and DeepSeek-Coder-Instruct-1.3B, 6.7B, and 33B [17]. These models exhibit varying levels of performance in code generation. We include more setup details about models in Appendix H.
Attention Steering Baseline. PASTA [62] is a recent method designed to steer model attention for better instruction following. Unlike Spa, PASTA requires a model-specific and time-consuming profiling stage to identify attention headers that are beneficial to performance. For each unique task, PASTA requires around 1000 training samples to identify attention headers that are beneficial to performance. By contrast, Spa is model-agnostic and only needs tuning a single hyperparameter with a few samples, which is fast and generalizable as detailed in Appendix J. PASTA internally edits transformers’ self-attention during the feed-forward process, whereas Spa only edits the final logits based on a mathematical approximation.
Prompting Methods. In addition to PASTA, we also compare Spa to the mainstream prompting-based code generation methods, including Self-Debugging [4], Self-Planning [21], ReAct [58], and Self-Edit [61]. Self-Debugging and Self-Edit leverage error messages from test cases to refine generated code. Self-planning generates a step-by-step plan before code generation. ReAct prompts an LLM to generate reasoning traces and action plans in an interleaved manner.
@@ 4.3 Evaluation Metrics
Following prior work [3], we measure code generation performance using the Pass@ metric, which measures whether any of the top candidates can pass all the test cases. For main results, we report Pass@ using greedy sampling to generate a single code snippet. To demonstrate the generalizability of Spa, we further calculate Pass@ using beam search in Appendix F.
@@ 4.4 Anchoring Strength Tuning
The anchoring strength serves as a hyperparameter in Spa. For each model and dataset, we use grid search to tune the anchoring strength on 1/5 of the tasks, and evaluate Pass@1 of Spa on the remaining 4/5 of the tasks. This process is repeated across all five folds, with final performance metrics averaged across folds.444We tune Spa on all tasks for better generalizability, while Spa is only activated for failed tasks during inference to optimize computational efficiency and generation performance. We observe an unimodal relationship between and the performance. We provide more details for the tuning Spa in Appendix J.
| Model | Size | HumanEval | HumanEval+ | MBPP | MBPP+ | BigCodeBench |
| CodeGen-Mono | (350M) | 15.3 | 12.2 | 19.6 | 15.9 | 1.1 |
| + Spa | 20.1 | 17.1 | 27.4 | 22.6 | 1.6 | |
| DeepSeek-Coder | (1.3B) | 66.4 | 61.8 | 58.2 | 52.4 | 2.5 |
| + Spa | 70.1 | 67.7 | 60.9 | 52.4 | 3.4 | |
| DeepSeek-Coder | (6.7B) | 75.6 | 70.2 | 67.0 | 58.5 | 12.7 |
| + Spa | 88.5 | 79.9 | 71.0 | 60.7 | 16.4 | |
| CodeLlama | (7B) | 33.6 | 28.2 | 50.9 | 40.8 | 3.4 |
| + Spa | 44.0 | 36.0 | 54.3 | 44.0 | 4.1 | |
| StarCoder2 | (16B) | 67.7 | 60.4 | 78.0 | 65.1 | 13.3 |
| + Spa | 75.6 | 65.6 | 82.0 | 69.1 | 14.3 | |
| DeepSeek-Coder | (33B) | 81.7 | 77.1 | 73.4 | 63.2 | 18.9 |
| + Spa | 86.2 | 79.3 | 79.4 | 70.3 | 21.5 |
@@ 5.1 Improvement over Base Models
Table 1 demonstrates that Spa consistently improves Pass@1 performance across all 5 benchmarks and 6 code LLMs. On HumanEval/HumanEval+ and MBPP/MBPP+, Spa enhances the base model with on average 5.5% absolute improvement and 14.5% relative improvement, achieving up to a 12.9% absolute improvement and a 42% relative improvement. These improvements are observed across models of varying sizes (350M-33B), original performance (15%-86%), and architectures. Notably, with Spa, the smaller DeepSeek-Coder (6.7B) outperforms its much larger 33B counterpart on HumanEval. This implies that optimizing model attention could provide a promising alternative to simply scaling up model size [22] for improving performance. On BigCodeBench, while the absolute improvements (average 1.56%) are more modest compared to other benchmarks, the relative gains (average 22.83%) remain significant. This is because Spa leverages the ability of its base model, if the base model could solve a task but overlooks a few important tokens in the prompt, Spa can help with this by adjusting the attention. If a model lacks the ability to solve a task, adjusting the model attention will not help much.
To demonstrate the generalizability of Spa, we show that Spa can consistently improve Pass@10 by, on average 3.42% and up to 7.9%, as detailed in Appendix F. We further evaluate Spa in scenarios where test cases are not available and show that it still significantly enhances performance, achieving an average 4.7% Pass@1 improvement on HumanEval, as detailed in Appendix K. To better investigate how Spa effectively anchors LLM’s attention on the user prompt, we provide two code generation examples in Appendix D. Spa achieves significant performance by only aligning model attention more closely with user prompts, without fine-tuning parameters or introducing additional context. We believe this success comes from its ability to improve attention reliability. We provide a thorough discussion in Appendix G.
@@ 5.2 Comparison to SOTA Methods
PASTA. Table 2 shows that Spa outperforms PASTA by achieving a 5.9X higher Pass@1 improvement while using only 20% of the inference time.555For a fair comparison, we average and add the latency of PASTA’s model profiling and the Spa’s tuning to the inference time for each task. We include more experimental details and discussion in Appendix L. Compared to the base model, Spa increases the decoding time by a factor of 1.27. We find that this latency is negligible in practice. Further discussion about the computational cost of Spa can be found in Appendix E.
@@ 5.3 Evaluation on Different Programming Languages
To evaluate the generalizability across different programming languages, we further evaluate Spa using HumanEval-X [18], which includes five programming languages.666In the latest version, Rust caused issues when running test cases, so we excluded Rust from the results. Table 3 demonstrates that Spa consistently improves Pass@1 on HumanEval-X, with an average increase of 7.9% for Python, 4.85% for Java, 6.5% for JavaScript, 3.65% for C++, and 5.2% for Go.

### v5

@@ 1 Introduction
Large language models (LLMs) have emerged as powerful programming assistants. They have demonstrated unprecedented capabilities in interpreting natural language descriptions of programming tasks and generating source code (Guo et al., 2024; Rozière et al., 2024). Despite this great progress, LLMs still produce incorrect solutions to challenging tasks or generate code that does not fully meet user expectations (Wang et al., 2024; Ning et al., 2023).
To improve the performance of LLMs on coding tasks, many efforts have been made to develop high-quality training data (Li et al., 2023; Guo et al., 2024; Wei et al., 2023) and design new domain-specific training objectives (Niu et al., 2022; Chakraborty et al., 2022). However, these approaches require tremendous computational resources. To address this challenge, various prompting methods have been explored to enhance the inference process of code LLMs, e.g., retrieval-augmented generation (Du et al., 2024), chain-of-thoughts (Le et al., 2024; Suzgun et al., 2022), self-planning and debugging (Jiang et al., 2024; Chen et al., 2024), etc.
@@ 2 Attention Analysis of Code LLMs
We conduct an empirical study to investigate the attention dilution phenomenon in code LLMs. Following prior studies (Zhang et al., 2022; Galassi et al., 2021), we obtain self-attention scores from the last layer in LLMs, which has been shown to represent the most accurate attention distribution (Kou et al., 2024; Wan et al., 2022a).111Intuitively, deeper layers capture representations with long-distance dependencies such as the control flow, which mirrors how humans understand programs. We calculate the percentage of attention on the user prompt. Calculation details are provided in Appendix B.4. We also experimented with an alternative gradient-based attention calculation method (Selvaraju et al., 2016) and obtained similar results, as detailed in Appendix B.2.
| Easy | Medium | Hard | Overall | |
| Passed | 294 | 475 | 400 | 390 |
| Failed | 418 | 664 | 784 | 622 |
We analyzed five code LLMs on HumanEval (Chen et al., 2021) and LiveCodeBench (Jain et al., 2024). Figure 2 shows the shift of LLMs’ attention on the user prompt during code generation. The result shows that as more tokens are generated, models’ attention on the user prompt gradually becomes less. Consequently, the code generation process becomes increasingly influenced by tokens generated in recent time steps, rather than the user prompt. This can be problematic in two ways. First, any generation inaccuracy, such as creating an unhelpful variable in earlier steps, is likely to propagate and influence subsequent steps. Second, when generating long and complex code with many detailed requirements, the model may overlook some specifications in the early stages, as the model tends to focus more on recent tokens. To further investigate whether this phenomenon affects performance, we experimented on LiveCodeBench (Jain et al., 2024), which provides 3 difficulty levels for each task. Table 1 shows that, for tasks with the same difficulty level, the average length of incorrectly generated code is consistently longer than that of correctly generated code. This implies that more generation steps can indeed cause attention dilution and hinder accuracy. More details are included in the Appendix A.
@@ 3.1 Overview
We propose Selective Prompt Anchoring (Spa) to steer model attention by amplifying the contextual impact of the important tokens in the user prompt. In this work, we assume these important tokens are selected by users. Inspired by the anchoring effect (Furnham & Boo, 2011) in psychology, we call these important tokens “anchored text”. For example, in the prompt, “Given a string, count the number of uppercase vowels”, the user may want to emphasize “uppercase vowels” to ensure that the model does not forget to define uppercase vowels (i.e., “A, E, I, O, U”) in the generated code.
@@ 3.2 Attention Steering and Prompt Anchoring
Spa performs attention steering by scaling the impact of selected tokens to the output logits . Suppose is the anchored text for which we want to adjust the impact. is an input embedding matrix at step , and represents a submatrix within corresponding to the . They are visualized below:
@@
To construct , we only retain columns of that correspond to , while setting remaining columns to zero. This matrix remains constant during token generation.
Conversely, is formed by zeroing out the same columns of that correspond to , while preserving all elements in the remaining columns.
@@
encapsulates the semantics for the anchored text , while encapsulates the semantics for the remaining text, such as the generated code. The sum of and reconstructs the original matrix :
Suppose we want to scale the semantic impact of the matrix by a value .222Scaling the semantics of by is not equivalent to multiplying by . Multiplying the embedding of by does not simply improve the “semantic influence” of , since the embedding of also encodes other non-semantic information such as positional information. This is why we compute the difference between the logits when masking and unmasking to cancel out noise and some of the non-semantic information (detailed in Section 3.3). We refer to as anchoring strength. We use to represent the function that augments by scaling the influence of by a degree of in the final logit.
@@
In this work, we focus on scaling up the impact of the anchored text to mitigate the attention dilution issue. Let represent the augmented logits calculated by model at step , where the impact of anchored text is scaled by . We can calculate the integral of the partial derivative of with respect to from 0 to .333 is differentiable for backpropagation. Formally,
@@ 3.3 Augmented Logits by Approximation
where the the integral part in Equation 7 is approximated by . To calculate , 444 does not mean setting the embedding vector to zeros. Instead, it means setting to zero, which replaces the original embedding for anchored text with the mask embedding that contains no semantic information. This mask embedding vector is non-zero. Conversely, a zero embedding does not necessarily indicate the absence of semantic information. we mask the anchored text using mask embeddings. Each LLM provides special tokens reserved for text masking, which almost has no semantic impact, e.g., <unk> for Code Llama (Rozière et al., 2024) and <pad> for DeepSeek-Coder (Guo et al., 2024). Each special token corresponds to a mask embedding. By replacing embeddings of with mask embeddings, we get a masked input matrix . It ablates the semantic impact of the anchored text while ensuring that the positional encoding remains unaffected. Thus, we can get
@@
We choose not to directly modify self-attention layers to steer the model attention, since this requires identifying which attention head in which layer to steer and the direct editing of attention values does not synchronize with other components like the feedforward layers, which can be fairly brittle and costly. Instead, we chose to simulate attention steering by logits manipulation, which is fast, reliable, and model-agnostic. In the next section, we demonstrate that Spa achieves better performance with less computational overhead compared to a SOTA method that directly modifies self-attention layers (Zhang et al., 2024).
4 Experiment Setup
@@ 4.1 Benchmarks
HumanEval (Chen et al., 2021) includes 164 Python tasks designed by OpenAI and now has become a widely-used benchmark for code generation.
MBPP (Austin et al., 2021) is another popular benchmark that includes 974 crowd-sourced Python tasks. Due to ambiguous task descriptions, the authors of MBPP created a sanitized version that included 427 tasks with clearer descriptions. We evaluate Spa on the sanitized version.
HumanEval+ and MBPP+ (Liu et al., 2023) improves the original HumanEval and MBPP benchmarks with additional test cases to cover corner cases (Liu et al., 2024b).
HumanEval-X (Hendrycks et al., 2021a) extends the HumanEval benchmark to support more programming languages such as Python, Java, JavaScript, C++, and Go. It aims to evaluate the multilingual code generation abilities.
BigCodeBench (Zhuo et al., 2024) is a more challenging benchmark for code generation that evaluates models’ abilities to follow complex instructions and use tools, including 1,140 real-world Python tasks across 139 libraries.
LiveCodeBench (Jain et al., 2024) is a contamination-free benchmark sourced from competitive programming platforms. The benchmark continues to evolve and add new code generation tasks. We conducted experiments on the latest release, which includes a total of 1,055 tasks. The latest tasks span from October 1, 2024, to February 1, 2025.
@@ 4.2 Baselines
Base Models. We select six representative open-source code LLMs: CodeGen-Mono-350M (Nijkamp et al., 2023), CodeLlama-7B (Rozière et al., 2024), StarCoder2-15B (Lozhkov et al., 2024), and DeepSeek-Coder-Instruct-1.3B, 6.7B, and 33B (Guo et al., 2024). These models exhibit varying levels of performance in code generation. We include more setup details about models in Appendix H.
Attention Steering Baseline. PASTA (Zhang et al., 2024) is a recent method designed to steer model attention for better instruction following. Unlike Spa, PASTA requires a model-specific and time-consuming profiling stage to identify attention headers that are beneficial to performance. For each unique task, PASTA requires around 1000 training samples to identify attention headers that are beneficial to performance. By contrast, Spa is model-agnostic and only needs tuning a single hyperparameter with a few samples, which is fast and generalizable as detailed in Appendix J. PASTA internally edits transformers’ self-attention during the feed-forward process, whereas Spa only edits the final logits based on a mathematical approximation.
Prompting Methods. In addition to PASTA, we also compare Spa to the mainstream prompting-based code generation methods, including Self-Debugging (Chen et al., 2024), Self-Planning (Jiang et al., 2024), ReAct (Yao et al., 2023), and Self-Edit (Zhang et al., 2023b). Self-Debugging and Self-Edit leverage error messages from test cases to refine generated code. Self-planning generates a step-by-step plan before code generation. ReAct prompts an LLM to generate reasoning traces and action plans in an interleaved manner.
@@ 4.3 Evaluation Metrics
Following prior work (Chen et al., 2021), we measure code generation performance using the Pass@ metric, which measures whether any of the top candidates can pass all the test cases. For main results, we report Pass@ using greedy sampling to generate a single code snippet. To demonstrate the generalizability of Spa, we further calculate Pass@ using beam search in Appendix F.
@@ 4.4 Anchoring Strength Tuning
The anchoring strength serves as a hyperparameter in Spa. For each model and dataset, we use grid search to tune the anchoring strength on 1/5 of the tasks, and evaluate Pass@1 of Spa on the remaining 4/5 of the tasks. This process is repeated across all five folds, with final performance metrics averaged across folds.555We tune Spa on all tasks for better generalizability, while Spa is only activated for failed tasks during inference to optimize computational efficiency and generation performance. We observe an unimodal relationship between and the performance. We provide more details for the tuning Spa in Appendix J.
@@ 5 Results
| Model | Size | HumanEval | HumanEval+ | MBPP | MBPP+ | BigCodeBench | LiveCodeBench |
| CodeGen-Mono | (350M) | 15.3 | 12.2 | 19.6 | 15.9 | 1.1 | 1.1 |
| + Spa | 20.1 | 17.1 | 27.4 | 22.6 | 1.6 | 1.5 | |
| DeepSeek-Coder | (1.3B) | 66.4 | 61.8 | 58.2 | 52.4 | 2.5 | 6.5 |
| + Spa | 70.1 | 67.7 | 60.9 | 52.4 | 3.4 | 9.2 | |
| DeepSeek-Coder | (6.7B) | 75.6 | 70.2 | 67.0 | 58.5 | 12.7 | 7.8 |
| + Spa | 88.5 | 79.9 | 71.0 | 60.7 | 16.4 | 10.8 | |
| CodeLlama | (7B) | 33.6 | 28.2 | 50.9 | 40.8 | 3.4 | 3.8 |
| + Spa | 44.0 | 36.0 | 54.3 | 44.0 | 4.1 | 4.0 | |
| StarCoder2 | (16B) | 67.7 | 60.4 | 78.0 | 65.1 | 13.3 | 7.0 |
| + Spa | 75.6 | 65.6 | 82.0 | 69.1 | 14.3 | 8.2 | |
| DeepSeek-Coder | (33B) | 81.7 | 77.1 | 73.4 | 63.2 | 18.9 | 11.9 |
| + Spa | 86.2 | 79.3 | 79.4 | 70.3 | 21.5 | 15.8 |
@@ 5.1 Improvement over Base Models
Table 2 shows Spa consistently improves Pass@1 across all 6 benchmarks and 6 code LLMs. On HumanEval/HumanEval+ and MBPP/MBPP+, Spa enhances the base model with on average 5.5% absolute improvement and 14.5% relative improvement, achieving up to a 12.9% absolute improvement and a 42% relative improvement. These improvements are observed across models of varying sizes (350M-33B), original performance (15%-86%), and architectures. Notably, with Spa, the smaller DeepSeek-Coder (6.7B) outperforms its much larger 33B counterpart on HumanEval. This suggests optimizing model attention is a promising alternative to simply scaling up model size (Kaplan et al., 2020) for performance improvement. On BigCodeBench and LiveCodeBench, while the absolute improvements (average 1.57% and 1.90%) are more modest compared to other benchmarks, the relative gains (average 22.83% and 28.50%) remain significant. This is because Spa leverages the ability of its base model. If the base model could solve a task but overlooks a few important tokens in the prompt, Spa can help with this by adjusting the attention. If a model lacks the ability to solve a task, adjusting the model attention will not help much.
To demonstrate the generalizability of Spa, we show that Spa can consistently improve Pass@10 by, on average 3.42% and up to 7.9%, as detailed in Appendix F. We further evaluate Spa in scenarios where test cases are not available and show that it still significantly enhances performance, achieving an average 4.7% Pass@1 improvement on HumanEval, as detailed in Appendix K. We illustrate Spa’s attention anchoring with two concrete examples in Appendix D. Spa improves performance by simply aligning attention to prompts, without introducing additional model parameters or context. We believe this success comes from its ability to improve attention reliability. We provide a thorough discussion in Appendix G.
@@ 5.2 Comparison to SOTA Methods
PASTA. Table 3 shows that Spa outperforms PASTA by achieving a 5.9X higher Pass@1 improvement while using only 20% of the inference time.666For a fair comparison, we average and add the latency of PASTA’s model profiling and the Spa’s tuning to the inference time for each task. We include more experimental details and discussion in Appendix L. Compared to the base model, Spa increases the decoding time by a practically negligible factor of 1.27. Further discussion about the computational cost of Spa can be found in Appendix E.
@@ 5.3 Evaluation on Different Programming Languages
To evaluate the generalizability across different programming languages, we further evaluate Spa using HumanEval-X (Hendrycks et al., 2021a), which includes five programming languages.777In the latest version, Rust caused issues when running test cases, so we excluded Rust from the results. Table 4 demonstrates that Spa consistently improves Pass@1 on HumanEval-X, with an average increase of 7.9% for Python, 4.85% for Java, 6.5% for JavaScript, 3.65% for C++, and 5.2% for Go.